2 bar magnets A and B of equal length, A having thrice the pole strength of B , are placed at right angle to each other with their north poles in contact. The system is then placed on afloating piece of cork so as to move freely. Determine (a)the angle that the magnet A forms with the magnetic meridian ,(b)the angle through which the system will rotate when the weaker magnet B is reversed.(H=0.4 OESTED)  

Grade:11

1 Answers

ROSHAN MUJEEB
3 years ago

N2S1=N1S21+N2S22−−−−−−−−−−−√N2S1=N1S12+N2S22
=L2+L2−−−−−−−√=L2–√=L2+L2=L2
∴∴Magnetic moment of the system
M=M=pole strength×N2S1×N2S1
M=m.L2–√=2–√mLM=m.L2=2mL